an iron rod 1 meter long is to be cut into two parts so that 1/2 of one is equal to 1/3 of the other
Answer by TimothyLamb(4379)

x = length of longer piece
y = length of shorter piece
---
x + y = 1
y/2 = x/3
3y = 2x
---
put the system of linear equations into standard form
---
x + y = 1
2x - 3y = 0
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = 0.6 m
y = 0.4 m
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---